### v2

# Enhancing Target-unspecific Tasks through a Features Matrix

Abstract
Recent developments in prompt learning of large vision-language models have significantly improved performance in target-specific tasks. However, these prompt optimizing methods often struggle to tackle the target-unspecific or generalizable tasks effectively. It may be attributed to the fact that overfitting training causes the model to forget its general knowledge having strong promotion on target-unspecific tasks. To alleviate this issue, we propose a novel Features Matrix (FM) regularization approach designed to enhance these models on target-unspecific tasks. Our method extracts and leverages general knowledge, shaping a Features Matrix (FM). Specifically, the FM captures the semantics of diverse inputs from a deep and fine perspective, preserving essential general knowledge, which mitigates the risk of overfitting. Representative evaluations demonstrate that: 1) the FM is compatible with existing frameworks as a generic and flexible module, and 2) the FM significantly showcases its effectiveness in enhancing target-unspecific tasks, achieving state-of-the-art performance.
Machine Learning, ICML
[…]
Figure 1: Our method is orthogonal to representative architectures, such as MaPLe (M) and PromptSRC (S), surpassing the existing state-of-the-art easy-to-use DePT by a significant margin.
Table 1: Our method is an easy-to-use design, integrated into CoOp, obtaining a higher average performance (11 datasets) on novel (target-unspecific) classes.
Method Design Base Novel
CLIP (ICML2021) Hand 69.34 74.22
CoCoOp (CVPR2022) Tune 80.47 71.69
PLOT (ICLR2023) Tune 81.24 72.98
CoOp (IJCV2022) Tune 82.38 67.96
+ DePT (CVPR2024) Tune 83.66 71.82
+ Kg (CVPR2023) Tune + Hand 80.73 72.70
+ Ours Tune + Hand 81.15 74.66
1 Introduction
Large vision-language models such as CLIP  (Radford et al., 2021) have attracted increasing attention for remarkable generalization performance. Vision-language models (VLMs) are trained to align textual and visual modalities by leveraging extensive datasets. For instance, a prominent example of such models is CLIP, which has achieved remarkable success across a wide range of downstream tasks (Greer et al., 2024; Sanghi et al., 2021; Etchegaray et al., 2024). CLIP utilizes a large collection of 400 million text-image pairs. One of CLIP’s most appealing features is its ability to perform zero-shot inference. During inference, CLIP utilizes hand-crafted text inputs, known as prompts, to generate classification weights to predict image features, all without requiring any target-specific parameters training.
In contrast to hand-crafted prompts, a model-parameter tuning method has been proposed as prompt learning to automatically learn prompt embeddings (Zhou et al., 2022a). For instance, CoOp (Zhou et al., 2022b) represents the first method that specifically focuses on learning the text embeddings of prompts with few-shot samples training while keeping the CLIP model frozen. Although CoOp (Zhou et al., 2022b) has shown significant performance improvements over hand-crafted prompts in target-specific base classes, it may yield inferior performance compared to the hand-crafted prompt CLIP in generalization tasks, e.g., the novel class of generalization from base-to-novel (see Table 1).
To overcome this challenge, some effective methods (Lu et al., 2022; Zhou et al., 2022a, b; Yao et al., 2023a; Chen et al., 2022) with tuning textual embeddings have been proposed. These methods aim to enhance the performance of novel classes, surpassing the novel performance of the previous CoOp. Although the novel ability of these methods surpasses that of CoOp ($67.96\%$), it is unfortunate that the novel ability of these methods still falls short compared to the hand-crafted prompt method (CLIP: $74.22\%$), as shown in Table 1 (Novel). A possible reason (Zhou et al., 2022a) may be that these prompting methods with tuning text embeddings tend to overfit the downstream data distributions. This overfitting can result in the model losing its inherent generalization capabilities (Yao et al., 2023a) obtained from hand-crafted prompts (Radford et al., 2021).
To alleviate this problem, we propose a novel Features Matrix (FM) with CLIP for target-unspecific tasks. Our method incorporates multiple hand-crafted prompts to extract general information as a pre-trained matrix from frozen CLIP to enhance generalization. The matrix of pre-trained features can delve into the semantics of different hand-crafted prompts finely and deeply, which reduces the risk of forgetting the essential general knowledge of pre-trained CLIP. Importantly, our method is compatible with current prompt learning frameworks for textual or multi-modal prompt learning and serves as a flexible and generic module. As demonstrated in Figure 1, our method integrated to the representative Maple (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b), exhibits competitive results in base-to-novel generalization, domain generalization, and cross-dataset generalization across 11 datasets. Meanwhile, our method surpasses the existing state-of-the-art easy-to-use DePT (Zhang et al., 2024) by a significant margin. Our contributions can be summarized as follows:
  • The proposed method incorporates multiple hand-crafted prompts with classes to extract general knowledge as a pre-trained features matrix (FM).
  • We propose an easy-to-use design, which is compatible with representative textual or multi-modal prompt learning frameworks for adapting CLIP.
  • Various target-unspecific tasks (base-to-novel generalization, cross-dataset generalization, and domain generalization) across 11 datasets demonstrate that our method showcases its effectiveness.
Figure 2: Illustration of representative textual prompting frameworks (CoOp and CoCoOp) and multi-modal prompting frameworks (MaPLe and PromptSRC). We propose a flexible and generic design, which is compatible with these representative architectures. In the figure, ”snowflake pattern” represents parameter freezing, ”flame pattern” represents learnable pattern, ”Deep” represents learnable tokens embedded in several layers of the encoder, where ”T” represents learnable text embedding, ”V” represents learnable visual embedding, ”Class priors” represents which category it belongs to, ”Tuning Similarity” represents the calculation of cosine similarity for fine-tuning architecture, and the light gray ”Similarity” represents the calculation of cosine similarity for frozen architecture. In the MaPLe architecture diagram, the ”e” represents the matrix function that connects the encoders of two modalities in several layers.
2 Preliminaries
Notations. Considering a pre-trained VLM, let $\mathcal{E}_{v}(\cdot)$ be its image encoder and $\mathcal{E}_{t}(\cdot)$ be its text encoder. The image encoder transforms input images into feature embeddings, capturing the visual information within the images. The text encoder generates representations for word embedding sequences, capturing the semantic information conveyed by the text prompts $\mathbf{p}$. Generally, a hand-crafted prompt $\mathbf{p}$ may have the form of “a photo of a [Class]”. In this paper, $\mathbf{x}$ represents an arbitrary image, and $l$ denotes the label.
Hand-Crafted CLIP. During the pre-training phase of CLIP (Radford et al., 2021), the image and text encoders are jointly trained on large-scale text-image pairs using a contrastive loss. This loss maximizes the cosine similarity between matching pairs and minimizes it between non-matching pairs, enabling the encoders to learn aligned visual and textual representations effectively. The final prediction score between the image $\mathbf{x}$ and text prompt $\mathbf{p}$ is computed using contrastive learning. The final prediction probability of alignment is computed by the matching score as follows:
$p(l=k\mid\mathbf{x})=\frac{\exp\left\{\operatorname{cos}\left(\mathcal{E}_{t}\left(\mathbf{p}_{k}\right),\mathcal{E}_{v}(\mathbf{x})\right)/\tau\right\}}{\sum_{k^{\prime}=1}^{K}\exp\left\{\operatorname{cos}\left(\mathcal{E}_{t}\left(\mathbf{p}_{k^{\prime}}\right),\mathcal{E}_{v}(\mathbf{x})\right)/\tau\right\}},$ (1)
where $l$ is the label of $\mathbf{x}$, $\operatorname{cos}(\cdot,\cdot)$ stands for cosine similarity between two vectors, and $\tau>0$ represents a temperature parameter. Here, the classifier consists of $K$ textual features derived from prompts $\left\{\mathbf{p}_{k^{\prime}}\right\}_{k^{\prime}=1}^{C}$, where the prompt $\mathbf{p}_{k^{\prime}}$ for the $k^{\prime}$-th class may have the form of “a photo of a”.
Textual Prompting. In textual prompting, the class name is retained as prior knowledge, while the word embeddings (referred to as context) of prompts are treated as learnable parameters, as shown in CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) of Figure 2. By modeling these context embeddings as trainable parameters, the model can optimize the prompts based on the specific requirements, enhancing the alignment between images and prompts. For class $k$, the tuning feature of the text encoder is denoted as ${t}^{tun}_{k}$ in a dataset with total $K$ classes. The final prediction probability of cross-entropy loss for two-modalities alignment is computed as:
$p(l=k\mid\mathbf{x})=\frac{\exp\left\{\operatorname{cos}\left({t}^{tun}_{k},\mathcal{E}_{v}(\mathbf{x})\right)/\tau\right\}}{\sum_{k^{\prime}=1}^{K}\exp\left\{\operatorname{cos}\left({t}^{tun}_{k^{\prime}},\mathcal{E}_{v}(\mathbf{x})\right)/\tau\right\}},$ (2)
where $\operatorname{cos}(\cdot,\cdot)$ stands for cosine similarity, and $\tau$ represents a temperature parameter. In some learned text embedding methods, the method aligns with the CoOp methodology by setting the embeddings to be shared across different classes. During the inference phase, the prompts, along with the learned embeddings, can generate textual features that are used for classification purposes. By leveraging these learned embeddings, the model can produce more effective representations, leading to improved target-specific classification performance (Zhou et al., 2022b).
Multi-Modal Prompting. Based on textual deep prompting, this design uses only a limited number of trainable parameters in the image encoder, as shown in MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) of Figure 2. The visual prompts are introduced at every transformer layer’s input space. In the visual branch, the input image $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$ is divided into $M$ patches. And, a tuning embedding of class $i_{class}$ is appended with the input patches. The foundation of vision prompting is ViT (Vision Transformer) (Dosovitskiy et al., 2020) backbone, which shares the same image encoder as CLIP. Let $\mathcal{P}_{v}$ denote the visual embeddings of prompts, and the image encoder processes the input tokens to generate tuning visual features, as follows (Jia et al., 2022):
$\tilde{\mathbf{x}}_{p}=\left\{\mathcal{P}_{v},\boldsymbol{i}_{class},\boldsymbol{i}_{1},\boldsymbol{i}_{2},\cdots,\boldsymbol{i}_{M}\right\},$ (3)
the tuning visual embeddings are introduced in the image encoder for deep prompt tuning.
Figure 3: Illustration of our easy-to-use method. We propose a novel Features Matrix (FM) for enhancing target-unspecific tasks. Our method incorporates multiple hand-crafted prompts with classes to extract general knowledge as a pre-trained features matrix. Various generalization tasks across 11 datasets demonstrate that our method outperforms existing prompt learning methods.
3 Methodology
3.1 Challenge and Motivation
As shown in Table 1, such as textual prompt (Zhou et al., 2022a, b; Yao et al., 2023a; Lu et al., 2022) and multi-modal prompting (Chen et al., 2022), have shown significant improvements in target-specific base classes. However, these methods often suffer from overfitting, leading to poor generalization in novel classes. To improve the novel performance of models, prompting regularization involves constraining the pre-trained and fine-tuning features in the text branch through computed loss. This constraint helps prevent forgetting general knowledge and ensures that the model retains essential information for downstream tasks. The objective of this design is to activate and maintain the pre-trained general knowledge, leveraging its remarkable generalization abilities to mitigate performance degradation when dealing with target-unspecific classes.
Recently, KgCoOp (Yao et al., 2023a) incorporates the constraint with $L1$ loss on fine-tuning feature with the pre-trained feature for the textual branch, avoiding the loss of CLIP’s original generalization capability. However, the novel performance of KgCoOp ($72.70\%$) is still lower than CLIP ($74.22\%$), as shown in Table 1 (Novel). It may be attributed to the fact that this single hand-crafted prompting regularization with CLIP cannot fully release and excavate the diversity semantics for utilizing essential general knowledge (Khattak et al., 2023b).
3.2 Proposed Features Matrix (FM)
To tackle this challenge and problem, we propose a novel design called Features Matrix (FM) for enhancing target-unspecific tasks. Our method incorporates multiple hand-crafted prompts with classes of the same datasets to extract general information as a pre-trained features matrix from CLIP to enhance generalization. In this way, we can explore and excavate the semantic information brought by each hand-crafted prompt more finely and deeply. By specifically aligning these multiple pre-trained hand-crafted unexpected features and tuning image features, our method aims to enhance the model’s generalization and robustness. This involves focusing on indistinguishable features that pose a challenge to the model during training.
As shown in Figure 3, we employ a set of hand-crafted prompt templates (P1, P2, P3, P4, etc) as inputs in the text encoder of pre-trained frozen CLIP. These prompt templates such as “a photo of one”, “a photo of a ”, “a picture of a”, “a drawing of a”, etc. These prompt templates are from the Prompt Engineering (Radford et al., 2021). In a dataset containing $K$ classes, the input to the text encoder is the frozen embeddings of hand-crafted prompts, consisting of different classes. The output of the frozen text encoder can be viewed as a features matrix, as depicted in Figure 3.
Specifically, we let $t$ denote the sets of pre-trained text features extracted by a frozen text encoder through a set of hand-crafted prompts templates (P1, P2, P3, P4, etc) with classes. In a dataset with total $K$ classes, features belonging to the current label $k$ of image samples are referred to as designated features $t_{k}$, while features from other classes are considered non-designated features $t_{\hat{k}}$. Our method computes the matching scores with the cosine distance for each visual output feature ${v}^{tun}$ and designated pre-trained text features of the features matrix as $\operatorname{cos}({t}_{k},{v}^{tun})$. Similarly, we compute the matching scores between each visual output feature and the non-designated pre-trained text features, denoted as $\operatorname{cos}(t_{\hat{k}},{v}^{tun})$. In correspondence with the features matrix, these matching scores can be viewed as a scores matrix, as depicted in Figure 3. Subsequently, we identify the textual feature with lower scores (low-$\beta$) among the designated features, which is referred to as unexpected designated features set ${F}^{k}_{un}$. Similarly, we identify textual features with higher scores (top-$\beta$) among the non-designated features, which are referred to as unexpected non-designated feature sets ${F}^{\hat{k}}_{un}$. Accordingly, the contrastive loss $\mathcal{L}_{\text{CL}}$ of unexpected features similarity is denoted as follows,
$\mathcal{L}_{\text{CL}}=-\log\frac{\exp\left\{\operatorname{cos}\left({t}_{k},{v}^{tun}\right)\right\}}{\exp\left\{\operatorname{cos}\left({t}_{k},{v}^{tun}\right)\right\}+\exp\left\{\operatorname{cos}\left({t_{\hat{k}}},{v}^{tun}\right)\right\}},$ (4)
where ${t}_{k}\in{F}^{k}_{un}$ and ${t_{\hat{k}}}\in{F}^{\hat{k}}_{un}$ as the selected unexpected features for designated unexpected features and non-designated unexpected features. The objective is to explore and excavate the unexpected features of general information. We train the objective for aligning the unexpected text features and tuning visual features, and the contrastive loss $\mathcal{L}_{\text{CL}}$ optimizes visual embeddings.
Table 2: Base-to-novel generalization. Our method achieves consistent average performance improvement over different baselines.
{floatrow}
(a) Average over 11 datasets
Base Novel HM
CoOp 82.69 63.22 71.66
+ Ours 81.15 74.66 77.79
Co 80.47 71.69 75.83
+ Ours 81.68 75.55 78.52
M 82.28 75.14 78.55
+ Ours 84.45 76.53 80.32
S 84.26 76.10 79.97
+ DePT 85.19 76.17 80.43
+ Ours 85.70 77.35 81.32
(b) ImageNet
Base Novel HM
CoOp 76.47 67.88 71.92
+ Ours 75.85 71.33 73.53
Co 75.98 70.43 73.10
+ Ours 77.35 72.36 74.79
M 76.66 70.54 73.47
+ Ours 78.18 71.38 74.62
S 77.60 70.73 74.01
+ DePT 78.20 70.27 74.02
+ Ours 78.90 71.58 75.07
(c) Caltech101
Base Novel HM
CoOp 98.00 89.81 93.73
+ Ours 97.58 96.60 97.13
Co 97.96 93.81 95.84
+ Ours 98.61 96.75 97.70
M 97.74 94.36 96.02
+ Ours 98.35 96.11 97.22
S 98.10 94.03 96.02
+ DePT 98.57 94.10 96.28
+ Ours 98.62 95.88 97.27
(d) OxfordPets
Base Novel HM
CoOp 93.67 95.29 94.47
+ Ours 93.78 97.80 95.78
Co 95.20 97.69 96.43
+ Ours 95.33 98.15 96.76
M 95.43 97.76 96.58
+ Ours 95.85 98.22 97.04
S 95.33 97.30 96.30
+ DePT 95.43 97.33 96.37
+ Ours 95.95 97.92 96.95
{floatrow}
(e) EuroSAT
Base Novel HM
CoOp 92.19 54.74 68.69
+ Ours 88.35 65.33 75.13
Co 87.49 60.04 71.21
+ Ours 88.15 70.11 78.12
M 94.07 73.23 82.35
+ Ours 94.22 75.65 83.93
S 92.90 73.90 82.32
+ DePT 92.23 77.90 84.88
+ Ours 95.50 76.85 85.17
(f) UCF101
Base Novel HM
CoOp 84.69 56.05 67.46
+ Ours 83.10 78.85 80.94
Co 82.33 73.45 77.64
+ Ours 82.95 77.21 80.00
M 83.00 78.66 80.77
+ Ours 87.33 79.10 83.02
S 87.10 78.80 82.74
+ DePT 87.73 77.70 82.46
+ Ours 88.80 79.50 83.93
(g) StanfordCars
Base Novel HM
CoOp 78.12 60.40 68.13
+ Ours 74.32 76.87 75.58
Co 70.49 73.59 72.01
+ Ours 72.88 76.10 74.46
M 72.94 74.00 73.47
+ Ours 78.66 75.13 76.86
S 78.27 74.97 76.58
+ DePT 80.80 75.00 77.79
+ Ours 80.91 76.51 78.68
(h) Flowers102
Base Novel HM
CoOp 97.60 59.67 74.06
+ Ours 96.22 72.32 82.57
Co 94.87 71.75 81.71
+ Ours 95.61 74.93 84.03
M 95.92 72.46 82.56
+ Ours 98.21 75.00 85.07
S 98.07 76.50 85.95
+ DePT 98.40 77.10 86.46
+ Ours 98.81 78.10 87.26
{floatrow}
(i) Food101
Base Novel HM
CoOp 88.33 82.26 85.19
+ Ours 89.98 92.85 91.41
Co 90.70 91.29 90.99
+ Ours 90.61 91.93 91.28
M 90.71 92.05 91.38
+ Ours 90.31 92.81 91.57
S 90.67 91.53 91.10
+ DePT 90.87 91.57 91.22
+ Ours 90.61 92.30 91.45
(j) FGVCAircraft
Base Novel HM
CoOp 40.44 22.30 28.75
+ Ours 37.32 34.61 35.92
Co 33.41 23.71 27.74
+ Ours 37.87 34.91 36.33
M 37.44 35.61 36.50
+ Ours 42.46 37.62 39.89
S 42.73 37.87 40.15
+ DePT 45.70 36.73 40.73
+ Ours 45.81 39.11 42.20
(k) SUN397
Base Novel HM
CoOp 80.60 65.89 72.51
+ Ours 79.12 78.38 78.77
Co 79.74 76.86 78.27
+ Ours 80.32 79.00 79.68
M 80.82 78.70 79.75
+ Ours 82.35 79.81 81.07
S 82.67 78.47 80.52
+ DePT 83.27 78.97 81.06
+ Ours 83.90 80.51 82.20
(l) DTD
Base Novel HM
CoOp 79.44 41.18 54.24
+ Ours 77.10 56.38 65.15
Co 77.01 56.00 64.85
+ Ours 78.90 59.61 67.93
M 80.36 59.18 68.16
+ Ours 83.01 60.98 70.32
S 83.37 62.97 71.75
+ DePT 84.80 61.20 71.09
+ Ours 84.90 62.58 72.07
For two-modal alignment, let $\mathcal{L}_{\mathrm{CE}}(\cdot,\cdot)$ denote the cross-entropy loss with prompts $\mathbf{p}$ for samples $\mathcal{S}$, as follows,
$\mathcal{L}_{\mathrm{CE}}=\arg\min_{\mathbf{p}}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{S}}\mathcal{L}\left\{\text{cos}\left({t}^{tun},{v}^{tun}\right),l\right\}.$ (5)
Accordingly, the objective $\mathcal{L}_{\text{total }}$ with a hyper-parameter $\gamma$ of our method can be formulated as follows,
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\gamma\mathcal{L}_{\text{CL}}.$ (6)
Consequently, optimizing models with $\mathcal{L}_{\text{CL}}$ between the pre-trained textual unexpected features and image tuning features. The $\mathcal{L}_{\text{CE}}$ represents two modalities of alignment of deep vision-language prompting. Importantly, our method is compatible with representative prompt learning frameworks as a generic and flexible module for textual prompt learning, such as CoOp and CoCoOp, or multi-modal prompt learning such as MaPLe and PromptSRC.
4 Main Generalization Tasks
4.1 Benchmark Setting
Table 3: Domain generalization.
Source Target
ImageNet -V2 -S -A -R Average
CoOp 71.51 64.2 47.99 49.71 75.21 59.28
+ DePT 72.63 64.80 48.05 50.00 75.50 59.58
+ Ours 71.82 65.13 48.10 50.15 76.13 59.87
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
+ DePT 72.77 65.10 49.10 51.00 76.85 60.51
+ Ours 72.10 65.25 50.13 52.11 77.18 61.16
MaPLe 70.72 64.07 49.15 50.9 76.98 60.27
+ DePT 73.27 65.33 49.05 51.25 77.50 60.78
+ Ours 71.56 65.45 50.33 52.32 78.10 61.55
PromptSRC 71.27 64.35 49.55 50.90 77.80 60.65
+ DePT 71.60 64.51 50.15 51.88 77.18 60.93
+ Ours 71.41 65.50 51.33 52.00 78.85 61.92
FAP 72.00 64.40 49.55 51.00 77.00 60.48
Table 4: Cross-dataset generalization. Our method achieves overall favorable performance.
Source Target
ImageNet Caltech101 OxfordPets StanfordCars Flowers102 Food101 Aircraft SUN397 DTD EuroSAT UCF101 Average
CoOp (Zhou et al., 2022b) 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
+ DePT 72.63 93.30 90.00 65.53 70.50 85.97 21.90 66.07 43.17 44.97 68.80 65.02
+ Ours 71.82 94.10 90.33 65.82 70.01 86.10 20.71 65.11 43.98 46.55 68.00 65.07
CoCoOp (Zhou et al., 2022a) 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
+ DePT 72.77 94.10 90.63 66.23 72.17 86.27 22.90 67.30 45.50 44.17 69.53 65.88
+ Ours 72.10 94.88 90.57 65.80 72.15 87.00 22.78 68.12 45.98 46.15 69.10 66.25
MaPLe (Khattak et al., 2023a) 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
+ DePT 73.27 92.53 90.10 64.60 70.10 85.57 23.63 66.40 45.03 40.13 67.53 64.56
+ Ours 71.56 94.00 90.91 65.92 73.10 86.78 24.33 68.35 46.13 49.01 68.81 66.73
PromptSRC (Khattak et al., 2023b) 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
+ DePT 71.60 93.80 90.13 66.00 70.93 86.27 24.30 67.23 46.60 45.83 69.10 66.02
+ Ours 71.41 94.95 92.14 66.50 73.25 87.53 25.81 68.10 49.01 49.20 69.74 67.62
FAP (Zhou et al., 2024) 72.00 93.40 90.15 65.61 70.13 85.89 23.70 67.88 45.98 45.12 68.99 65.68
Compared Methods. In three generalization experiments, we compare our method with CLIP (ICML2021) (Radford et al., 2021), CoOp (IJCV2022) (Zhou et al., 2022b), CoCoOp (CVPR2022) (Zhou et al., 2022a), MaPLe (M) (CVPR2023) (Khattak et al., 2023a), PromptSRC (S) (ICCV2023) (Khattak et al., 2023b), FAP (NIPS2024) (Zhou et al., 2024) and existing DePT (CVPR2024) (Zhang et al., 2024). The DePT is an easy-to-use method. We process frozen CLIP using a linear probe method. Comparison methods are based on ViT-B/16 architecture for fair comparison.
Implementation Details. We employ a CLIP model based on the ViT-B/16 architecture. For the PromptSRC-based and MaPLe-based frameworks, we set the visual and textual embedding length to 4. We set the easy-to-use module $\gamma$ to $0.1$ and matching scores (top and low) $\beta$ to $5$. Training for 30 epochs for a base-to-novel setting in the first 9 transformer layers. Training for 20 epochs for domain generalization setting and cross-dataset evaluation setting in the first 3 transformer layers. We use an SGD optimizer with a learning rate of 0.0025 on a single GPU. Following the hand-crafted CLIP, we use the hand-crafted template sets from (Radford et al., 2021).
4.2 Datasets
In Table 5, the datasets cover multiple recognition tasks. These datasets are all open-source datasets. We conducted base-to-novel generalization experiments, few-shot learning experiments, and cross-dataset generalization experiments on 11 datasets. We conduct domain generalization experiments on four variant datasets.
Dataset Classes Train Val Test
ImageNet (Deng et al., 2009) 1,000 1.28 M N/A 50,000
Caltech101 (Fei-Fei et al., 2004) 100 4,128 1,649 2,465
OxfordPets (Parkhi et al., 2012) 37 2,944 736 3,669
StanfordCars (Krause et al., 2013) 196 6,509 1,635 8,041
Flowers102 (Nilsback & Zisserman, 2008) 102 4,093 1,633 2,463
Food101 (Bossard et al., 2014) 101 50,500 20,200 30,300
FGVCAircraft (Maji et al., 2013) 100 3,334 3,333 3,333
SUN397 (Xiao et al., 2010) 397 15,880 3,970 19,850
DTD (Cimpoi et al., 2014) 47 2,820 1,128 1,692
EuroSAT (Helber et al., 2019) 10 13,500 5,400 8,100
UCF101 (Soomro et al., 2012) 101 7,639 1,898 3,783
-V2 (Recht et al., 2019) 1,000 N/A N/A 10,000
-Sketch (Wang et al., 2019) 1,000 N/A N/A 50,889
-A (Hendrycks et al., 2021b) 200 N/A N/A 7,500
-R (Hendrycks et al., 2021a) 200 N/A N/A 30,000
Table 5: Training and testing datasets. dataset consists of 11 image classification datasets and four variant datasets of ImageNet.
4.3 Base-to-Novel Generalization Task
In the base-to-novel generalization task, the datasets are divided into base and novel classes. The model is trained on the base classes in a 16-shot setting, and tested on both the base and novel classes across 11 different datasets. The number of classes for base and novel is the same, which means that all classes in a dataset are evenly divided into two groups of classes. The process of dividing all classes in the dataset is randomly selected. HM refers to harmonic mean. This experiment evaluates the generalization ability of methods. In Table 2, our method demonstrates significant improvements on all $11$ datasets on HM. Our method significantly enhances the performance of base classes while also improving the accuracy of zero-shot CLIP on novel classes. In terms of average performance, our method based on PromptSRC (s) attains $85.70\%$ accuracy on the base classes, and $77.35\%$ accuracy on the novel classes. Our method, as an easy-to-use module, achieves performance improvements for various baseline frameworks, such as CoOp, CoCoOp, MaPLe (M), and PromptSRC (S).
4.4 Cross-Dataset Generalization Task
We train our model with 16 shots on ImageNet and test on 10 other unseen datasets. This experiment aims to validate the potential of our method in a wide range of cross-dataset transfers. Our method for training ImageNet is lower than that of DePT. As shown in Table 4, our method based on PromptSRC shows competitive performance in $/10$ 10 over the generic and flexible method DePT. These findings suggest that our method excels in achieving better generalization across a diverse range of unseen datasets.
4.5 Domain Generalization Task
We train our model using the ImageNet in 16 shots and test its performance on 4 different variants of the ImageNet. In Table 3, our method based on PromptSRC demonstrates the highest average improvement of $1.27\%$ over the PromptSRC method. Additionally, when compared to DePT, our method consistently outperforms it across all ImageNet variant datasets. This indicates that our method is designed to enhance domain generalization.
5 Further Studies
5.1 Learning Depth and Embeddings Length
In Table 6, we note that increasing the learning depth generally increases the performance based on PromptSRC. As the number of layers increases to $11$, the HM value decreases. It indicates excessive fine-tuning of the model, causing it to lose CLIP generalization. In Table 7, our findings indicate that the performance based on PromptSRC reaches its peak when the length of embeddings is set to $4$ on HM for an average of 11 datasets. Our ablation studies are based on keeping all other settings unchanged.
Table 6: Analysis of learning depth based on PromptSRC for an average of 11 datasets.
Learning Depth 1 3 5 7 9 11
HM 77.11 78.03 79.12 80.01 81.32 80.21
Table 7: Analysis of embeddings length based on PromptSRC for an average of 11 datasets.
Prompt Length 1 2 4 6 8 10
HM 77.01 80.05 81.32 80.30 79.11 78.54
5.2 Analysis of $\gamma$ for $\mathcal{L}_{\text{CL}}$
In Table 8, it is observed that as the number of $\gamma$ to 0.1, there is a peak in HM for an average of 11 datasets. We conduct analysis based on the PromptSRC model.
5.3 Analysis of Top-$\beta$ and Low-$\beta$
In Table 9, it is observed that as the number of Top-$\beta$ and Low-$\beta$ to 5, there is a peak in HM ($81.32\%$). We conduct analysis based on the PromptSRC model for an average of 11 datasets. We keep other hyperparameters unchanged.
Table 8: Analysis of $\gamma$ for $\mathcal{L}_{\text{CL}}$ in pre-trained features matrix. We conduct analysis based on the PromptSRC (Khattak et al., 2023b) model for an average of 11 datasets. HM refers to harmonic mean.
Average (11 datasets)
0.05 0.1 0.5 0.9
HM ($\gamma$) 80.13 81.32 81.00 79.88
Table 9: Identifying scores of Top-$\beta$ and Low-$\beta$ for an average of 11 datasets. We conduct analysis based on the PromptSRC (Khattak et al., 2023b) model. HM refers to harmonic mean.
Average (11 datasets)
3 4 5 6
HM ($\beta$) 78.21 79.81 81.32 80.51
5.4 Applying to Other ViT Instances
In Table 10, we apply our method in various types of ViT instances. Our experimental performance is based on the average performance of 11 datasets. Our generic and flexible module based on PromptSRC has the lowest performance of ViT-B/32, with a value of $79.70\%$. And, our method, which leverages PromptSRC, achieves the best performance with ViT-L/14, with a value of $84.20\%$.
Table 10: Applying to other ViT instances.
Average (11 datasets)
ViT-B/32 ViT-B/16 ViT-L/14
HM (ViT) 79.70 81.32 84.20
5.5 Inference Stage Computational Cost
In Table 11, the compute cost analysis is performed using the SUN397 dataset over 10 epochs on a single GPU. While our method (row 3) may have a slower inference speed as a result of multiple cosine similarity calculations. The number of parameters we introduced has remained relatively consistent and stable in training. This implies that usually, our parameter quantity is lower than that of DePT.
Table 11: The compute cost analysis is performed using the SUN397 (Xiao et al., 2010) dataset over 10 epochs on a single GPU. ‘N’: the num of classes in the base task (Zhang et al., 2024).
Method Train time Learnable para. HM
CoOp 10.88min 8K 71.65
+ DePT 10.91min + (2+N/2)K 77.30
+ Ours 13.56min + 3072 78.66
6 Related Work
6.1 Vision Language Models
Recently, researchers have demonstrated the strong generalization capability of Visual Language Models (VLMs) (Alayrac et al., 2022) which involve training on large-scale datasets of image-text pairs. Such as CLIP (Radford et al., 2021), which is a prominent and straightforward framework among existing VLMs. The strong generalization capability of CLIP has made it a foundation for many methods in adapting pre-trained VLMs for downstream tasks (Sanghi et al., 2022; Maaz et al., 2022; Bangalath et al., 2022; Zhang et al., 2021). To enhance the generalization ability of VLMs, researchers have explored various approaches. Some approaches involve enhancing the text encoder or the visual encoder (Vaswani et al., 2017). By improving the capabilities of the text encoder, the model can better capture the semantics and contextual information in the textual input. Similarly, enhancing the visual encoder allows the model to extract more informative and discriminative features from the visual input. Another approaches deeply fuse the visual and text knowledge within the model. Using a large number of images for training is indeed one of the most effective methods.
6.2 Prompt Tuning in Vision Language Models
Prompt tuning (Lester et al., 2021) is a commonly employed technique in the field of Natural Language Processing (NLP) for training on downstream tasks. Leveraging text prompts, which are instructions given to the language model component of VLMs, is a prevalent practice to improve task comprehension. Full fine-tuning and linear probes are two commonly employed approaches for adapting VLMs to downstream tasks. However, the constraints of both methods have prompted research into innovative techniques influenced by prompt tuning within the realm of Natural Language Processing. CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) fine-tuning the CLIP model for few-shot image recognition by optimizing a continuous set of embeddings within the textual branch. The image-conditional prompt utilized in CoCoOp significantly contributes to improving generalization to unseen classes. Moreover, some approaches (Yao et al., 2023b, a) constrain the proposed learnable prompts contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022a) ensures that the language model focuses on pertinent visual information when making predictions. In addition to prompt tuning, approach (Zhang et al., 2024) introduces a generic and flexible approach to align its vision and language representations.
7 Limitations and Conclusion
Due to our method incorporating multiple contrastive learning processes, the training speed is slower. Moreover, we plan to investigate the potential of our method in other downstream tasks (Sanghi et al., 2022).
In this paper, we propose a Features Matrix (FM) for vision-language models. Our method incorporates multiple hand-crafted prompts to extract general information as a pre-trained features matrix from CLIP to enhance generalization. The matrix of pre-trained features can delve into the semantics of different hand-crafted prompts at a more profound level, which reduces the risk of forgetting the essential general knowledge of pre-trained CLIP. Importantly, our method is compatible with representative prompt learning frameworks for textual or multi-modal prompts as a generic and flexible module.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. None of these points we feel must be specifically highlighted here.